OpenStudy (anonymous):

PLEASE HELP?! Enter in standard form the equation of the line passing through the given point and having the given slope. B(2, -5), m = -2/3

OpenStudy (anonymous):

y-y1=m(x-x1) plug the values

OpenStudy (anonymous):

y - (-5) = -2/3 (x-2)

OpenStudy (anonymous):